# Large Orders and Strong-Coupling Limit in Functional Renormalization

### v1

@@ I Introduction
Perturbative expansions are a work horse in theoretical physics. Most of them are not converging, but aysmptotic series [1, 2, 3, 4, 5]. The main strategy to obtain a series with a finite radius of convergence is to define its Borel transform by dividing its $n$-th series coefficient by $n!$. One then continues the latter and reconstructs the original function via an integral over this analytic continuation. The aim is to extend the range of applicability from small expansion parameters, where the series naively converges, to larger ones. Techniques using Padé-Borel resummation or conformal mappings are successful here [3, 4, 6, 7, 8, 5], and were employed for the $\epsilon$-expansion of perturbative RG [8, 6, 9]. In simpler examples, as the anharmonic oscillator [10], one can go further, and use resurgence [11, 12, 13] to reach finite couplings. Resurgence is strong in identifying the closest singularity of the Borel transform, and using this information to extend the domain of convergence. It is less effective to reach strong coupling.
In the renormalization context we wish to take a bare coupling to infinity; we know of few methods to address this beyond postulating the asymptotic behavior [14]. An additional problem arises when the microscopic set of couplings is itself a function, as in the functional renormalization group (FRG) treatment of disordered systems. Here we consider a specific model in dimension $d=0$, which is later derived from the field theory of disordered elastic manifolds (for a review, see [15]). We wish to answer the following fundamental questions: What is the large-order behavior of FRG? Is it Borel-summable? How can we study its strong-coupling limit? And how does universality arise?
@@ II Setting the stage
$\tilde{\phi}^{n}f(\phi)\right>=(\partial_{\phi})^{n}f(\phi)\Big{|}_{\phi=0}\quad\Rightarrow\quad\left<\tilde{\phi}^{n}\phi^{m}\right>_{0}=n!\,\delta_{n,m}.$ (2)
@@
${\cal Z}_{O(2)}^{\rm B}(t)$ has a branch cut starting at $t=-1/4$, and its analytic continuation is well defined for $t>0$. This allows one to obtain ${\cal Z}_{O(2)}(\lambda)$ via an inverse Borel transform
@@ III Resummation of a functional expansion
Let us proceed to a 0-dimensional model for functional RG,
${\cal Z}_{\rm FRG}(w,\lambda)=\int_{\tilde{\phi},\phi}{\mathrm{e}}^{-\tilde{\phi}(\phi-w)+\lambda\tilde{\phi}^{2}[\Delta(0)-\Delta(\phi)]}.$ (6)
We consider this a mathematical problem. We assume that $\Delta(\phi)$ is an analytic function, fast and monotonously decaying for $\phi\geq 0$, and that $\Delta(0)=1$. A good example is $\Delta(\phi)={\mathrm{e}}^{-\phi}$.
@@
${\cal Z}_{\rm FRG}(w,\lambda)$ ::=:= $\sum_{n=0}^{\infty}\lambda^{n}{\cal Z}_{\rm FRG}^{(n)}(w),$ (7)
${\cal Z}_{\rm FRG}^{(n)}(w)$ ::=:= $\frac{1}{n!}(\partial_{w})^{2n}\big{[}1-\Delta(w)\big{]}^{n}.$ (8)
@@
$\quad g_{w}(\phi)$ ::=:= $\frac{1-\Delta(w{+}\phi)}{\phi^{2}}.$ (10)
@@
Figure 2: Plot of $[g_{w}(\phi)/g_{w}(\phi_{\rm SP})^{n}$ for $w=0$, $n=100$, with real part in blue (solid) and imaginary part in red (dashed); $\phi=\phi_{\rm SP}+iy/\sqrt{n}$, as indicated by the red curve on Fig. 1. In green (dot-dashed) $\exp(-\frac{g_{w}(\phi)}{\phi^{2}g_{w}^{\prime\prime}(\phi)}y^{2}/2)|_{\phi=\phi_{\rm SP}}$, whose integration leads to Eq. (14).
@@
${{\cal Z}_{\rm FRG}^{(n)}(w)\simeq\frac{\Gamma(2n{+}1/2)}{\Gamma(n{+}1)\sqrt{\pi}}{\left[{g_{w}(\phi)}\right]^{n}}\sqrt{\frac{g_{w}(\phi)}{\phi^{2}g_{w}^{\prime\prime}(\phi)}}\Bigg{|}_{\phi=\phi_{\rm SP}}}.$ (14)
Several remarks are in order: First of all, the large-order behavior is asymptotic and its Borel transform exists, as $\Gamma(2n{+}1/2)/\Gamma(n{+}1)\simeq n!$. Second, for disordered systems $\Delta(\phi)=\Delta(|\phi|)$ thus the branch for $\phi<0$ is given by the branch for $\phi>0$. In contrast, the saddle point is at negative $\phi$, on the analytic continuation of the branch for positive $\phi$, outside its physically relevant domain. A numerical check for $n=100$ is shown in Fig. 2. The relative error for ${\cal Z}_{\rm FRG}^{(n)}(0)$ is $10^{-4}$, which can systematically be improved by further $1/n$ corrections. The latter are relevant for resurgence [23].
There are surprises in this procedure: when changing the microscopic disorder from $\Delta(\phi)={\mathrm{e}}^{-\phi}$ to $\Delta(\phi)={\mathrm{e}}^{-\phi-a\phi^{2}}$, there is a critical value $a_{\rm c}\approx 0.0649$ s.t. if $a>a_{\rm c}$ the saddle point disappears. We will see later that while the large-order behavior changes, the large-$\lambda$ limit can still be taken, and is independent of $a$, as long as $a>0$. We retain that, despite its widespread use, the information contained in the large-order behavior may be quite limited [23].
@@ IV Borel transform
where $\phi_{0}\leq 0<\phi_{1}$ are the two zeros of the denominator, and the sign inside the square root is reversed between Eqs. (16) and (17). For $w=0$, $\phi_{0}=0$. One could extend this integral from $\infty$- ∞ to $\infty$, if one keeps only the real part of the integrand. A numerical check of Eqs. (16) and (17) is presented in Fig. 4 of the appendix.
@@ VI Strong-coupling behavior
Integrating twice yields
@@
What is remarkable about Eq. (20) is that the final result, given in Eq. (22), is largely independent of the microscopic $\Delta(\phi)$. What we used is that $\Delta(\phi)$ is analytic, has a linear cusp at the origin, and decays quickly. The cusp is a technical requirement, necessary to transform the contour integral into a cut integral. We believe that this is more a technical constraint than a physical one: we could regularize the microscopic disorder in order to obtain a linear cusp, and then remove the regularization. We have studied this for $\Delta(\phi)={\mathrm{e}}^{-\phi^{2}}$. While we clearly see that convergence is non-uniform and slow, we have no indication that the process does not converge, or converges against a different fixed point. On the practical side, when applied to disordered systems, as the disorder usually lives on a grid, we can well approximate it by a function with a linear cusp.
@@ VII Field theory for disordered elastic systems
Here $\tilde{\phi}$ and $\phi$ are bosonic fields (complex numbers), while $\bar{\psi}_{i}$ and $\psi_{i}$ are Grassmann variables [37]. Going to dimension $d=0$, dropping the Grassmann fields, and rescaling $\tilde{\phi}\to\tilde{\phi}/m^{2}$ yields the integral (6), with
@@
(Again, $\Delta(0)=1$). Eq. (23) implies that $w\sim\sqrt{\lambda}=m^{-2}$, thus the scaling exponent of the field $\phi$, known as the roughness exponent $\zeta$, is
$\zeta=2.$ (27)
Let us rewrite these findings in terms of the action (VII). Retaining only the bosonic fields, we get for the model (6)-(8),
${\cal Z}_{\rm FRG}(w,\lambda)\equiv{\cal Z}^{{\cal S}}_{\rm bos}(w,\lambda):=\int_{\phi,\tilde{\phi}}{\mathrm{e}}^{-{\cal S}|_{\psi_{i}\to 0}}.$ (28)
By construction, the partition function of action (VII) over all bosonic and Grassmann fields is $1$,
@@
This function has a limit (we use Eq. (26))
$\tilde{\Delta}_{\rm Susy}(w):=\lim_{\lambda\to\infty}\Delta_{\rm Susy}(w\sqrt{\lambda},\lambda).$ (31)
Surprisingly, the functions defined in Eqs. (31) and (24) agree,
@@
Thus what we obtained for the simple model (6) also applies to the disordered system defined by the action (VII), order by order in perturbation theory.
@@ VIII Applications
Our results agree up to 1-loop order with that for disordered elastic manifolds in equilibrium and at depinning [17, 19, 18, 20, 21, 22]. Beyond that, amplitudes are different in the $\epsilon$-expansion, and there are additional anomalous terms which are hard to recuperate [23]. Though our model can formally be derived from a field theory in equilibrium, we do not believe $\Delta_{\rm FRG}(w)$ to be relevant for a specific physical situation, even though the predicted roughness exponent is equal to that of depinning, and the shape of $\Delta_{\rm FRG}(w)$ on Fig. 3 is between a driven particle and Sinai’s model, both relevant in $d=0$.
Given these caveats, we turn to the strength of our approach. Our model contains all ingredients of functional renormalization: it shows that the perturbative series is Borel summable, how the limit of strong coupling is reached, that it cannot be inferred from the large-order behavior, and how universality emerges. This poses a solid framework for the strong-coupling behavior in functional renormalization. We also saw that to define the path integral non-perturbatively, one needs to specify the integration contours, and possibly restrict variables to part of their physically allowed domains.
@@ Appendix A Field Theory
Building on the Susy formulation of [38], Ref. [31] introduces two physical copies located at $u_{1}$ and $u_{2}$, which are subject to confining potentials displaced by $w$, s.t. their difference $\phi:=u_{1}-u_{2}$ has expectation $\phi\right>=w$; its center of mass is $u:=(u_{1}+u_{2})/2$. The field theory, given in Eq. (36) of [31] reads
@@
After Eq. (20) we stated that the latter can be derived from Eq. (6) if $\tilde{\phi}\in i\mathbb{R}$ and $\phi>0$. We use the same prescription to pass from Eq. (34) to Eq. (35). This allows us to evaluate Eqs. (28)-(30), both perturbatively and non-peturbatively.
@@ Acknowledgements.
We are grateful to Andrei Fedorenko for stimulating discussions and many deep questions. We also profited from discussions with Costas Bachas and Edouard Brézin.

### v2

@@ I Introduction
Perturbative expansions are a work horse in theoretical physics. Most of them are not converging, but asymptotic series [1, 2, 3, 4, 5, 6]. The main strategy to obtain a series with a finite radius of convergence is to define its Borel transform by dividing its $n$-th series coefficient by $n!$. One then continues the latter and reconstructs the original function via an integral over this analytic continuation. The aim is to extend the range of applicability from small expansion parameters, where the series naively converges, to larger ones. Techniques using Padé-Borel resummation or conformal mappings are successful here [3, 4, 7, 8, 9, 5], and were employed for the $\epsilon$-expansion of perturbative RG [9, 7, 10, 6]. In simpler examples, as the anharmonic oscillator [11], one can go further, and use resurgence [12, 13, 14] to reach finite couplings. Borel resummation identifies singularities of the Borel transform, and using this information extends the domain of convergence. It is not effective in reaching strong coupling.
An additional problem arises when the microscopic set of couplings is itself a function, as in the functional renormalization group (FRG) treatment of disordered systems. In FRG, one uses a confining potential of strength $m^{2}$ to obtain the effective disorder in terms of the bare one, order by order in $\lambda=m^{d-4}$. Varying the FRG scale $m$ allows one to obtain an approximate solution in the limit of $m\to 0$, i.e. $\lambda\to\infty$.
Here we consider a specific model in dimension $d=0$, which is later derived from the field theory of disordered elastic manifolds (for a review, see [15]), in which we can take the limit of $\lambda\to\infty$ directly. We wish to answer the following fundamental questions: What is the large-order behavior of FRG? Is it Borel-summable? How can we study its strong-coupling limit? And how does universality arise?
@@ II Setting the stage
$\left<\tilde{\phi}^{n}f(\phi)\right>=(\partial_{\phi})^{n}f(\phi)\Big|_{\phi=0}\quad\Rightarrow\quad\left<\tilde{\phi}^{n}\phi^{m}\right>_{0}=n!\,\delta_{n,m}.$ (2)
@@
${\cal Z}_{O(2)}^{\rm B}(t)$ has a branch cut starting at t=1/4t=-1/4, and its analytic continuation is well defined for $t>0$. This allows one to obtain ${\cal Z}_{O(2)}(\lambda)$ via an inverse Borel transform
@@ III Resummation of a functional expansion
Let us proceed to a 0-dimensional model for functional RG (norm as in Eq. (1)),
${\cal Z}_{\rm FRG}(w,\lambda):=\int_{\tilde{\phi},\phi}{\mathrm{e}}^{-\tilde{\phi}(\phi-w)+\lambda\tilde{\phi}^{2}[\Delta(0)-\Delta(\phi)]}.$ (6)
At this stage, this is a mathematical problem; we show later its significance for depinning. We assume that $\Delta(\phi)$ is an analytic function, fast and monotonously decaying for $\phi\geq 0$, and that $\Delta(0)=1$. A good example is $\Delta(\phi)={\mathrm{e}}^{-\phi}$.
@@
${\cal Z}_{\rm FRG}(w,\lambda)$ $:=$ $\sum_{n=0}^{\infty}\lambda^{n}{\cal Z}_{\rm FRG}^{(n)}(w),$ (7)
${\cal Z}_{\rm FRG}^{(n)}(w)$ $:=$ $\frac{1}{n!}(\partial_{w})^{2n}\big[1-\Delta(w)\big]^{n}.$ (8)
@@
$\quad g_{w}(\phi)$ $:=$ $\frac{1-\Delta(w{+}\phi)}{\phi^{2}}.$ (10)
@@
Figure 2: Plot of $g_{w}(\phi)/g_{w}(\phi_{\rm SP})]^{n}$ for $w=0$, $n=100$, with real part in blue (solid) and imaginary part in red (dashed); $\phi=\phi_{\rm SP}+iy/\sqrt{n}$, as indicated by the red curve on Fig. 1. In green (dot-dashed) $\exp(-\frac{g_{w}^{\prime\prime}(\phi)}{g_{w}(\phi)}\frac{y^{2}}{2})|_{\phi=\phi_{\rm SP}}$, whose integration leads to Eq. (14).
@@
${\cal Z}_{\rm FRG}^{(n)}(w)\simeq\frac{\Gamma(2n{+}1/2)}{\Gamma(n{+}1)\sqrt{\pi}}{\left[{g_{w}(\phi)}\right]^{n}}\sqrt{\frac{g_{w}(\phi)}{g_{w}^{\prime\prime}(\phi)}}\Bigg|_{\phi=\phi_{\rm SP}}$}. (14)
The large-order behavior is asymptotic and its Borel transform exists, as $\Gamma(2n{+}1/2)/\Gamma(n{+}1)\simeq n!$. The saddle point is at negative $\phi$, on the analytic continuation of the branch for $\phi\geq 0$, outside its physically relevant domain. A numerical check for $n=100$ is shown in Fig. 2. The relative error for ${\cal Z}_{\rm FRG}^{(n)}(0)$ is $10^{-4}$, which can systematically be improved by further $1/n$ corrections. The latter are relevant for resurgence [23].
When changing the microscopic disorder from $\Delta(\phi)={\mathrm{e}}^{-\phi}$ to $\Delta(\phi)={\mathrm{e}}^{-\phi-a\phi^{2}}$, there is a critical $a_{\rm c}\approx 0.0649$ beyond which the real saddle point at $w=0$ disappears. It is replaced by an infinity of pairs of complex saddle points, corresponding to a more intricate resurgent structure [23].
@@ IV Borel transform
where $\phi_{0}\leq 0<\phi_{1}$ are the two zeros of the denominator, and the sign inside the square root is reversed between Eqs. (16) and (17). For $w=0$, $\phi_{0}=0$. One could extend this integral from -$\infty$ to $\infty$, if one keeps only the real part of the integrand. A numerical check of Eqs. (16) and (17) is presented in Fig. 4 of the appendix.
@@ VI Strong-coupling behavior
Integrating twice and using $\tilde{\Delta}_{\rm FRG}(\infty)=0$ yields
@@
What is remarkable about Eq. (20) is that the final result, given in Eq. (22), is largely independent of the microscopic $\Delta(\phi)$. What we used is that $\Delta(\phi)$ is analytic, has a linear cusp at the origin, and decays quickly. The cusp is a technical requirement, necessary to transform the contour integral into a cut integral. We believe that this is more a technical constraint than a physical one: we could regularize the microscopic disorder to obtain a linear cusp, and then remove the regularization. We have studied this for $\Delta(\phi)={\mathrm{e}}^{-\phi^{2}}$. While we clearly see that convergence is non-uniform and slow, we have no indication that the process does not converge, or converges against a different fixed point. On the practical side, when applied to disordered systems, as the disorder usually lives on a grid, we can well approximate it by a function with a linear cusp.
@@ VII Field theory for disordered elastic systems
Here $\tilde{\phi}$ and $\phi$ are bosonic fields (complex numbers), while $\bar{\psi}_{i}$ and $\psi_{i}$ are Grassmann variables [37]. To understand this action, let us temporarily drop the fermionic fields. Taking dimension $d\to 0$, and rescaling $\tilde{\phi}\to\tilde{\phi}/m^{2}$, we get the model of Eqs. (6)-(8),
${\cal Z}_{\rm FRG}(w,\lambda)$ $\equiv$ ${\cal Z}^{{\cal S}}_{\rm bos}(w,\lambda)\big|_{d=0}:=\int_{\phi,\tilde{\phi}}{\mathrm{e}}^{-{\cal S}|_{\psi_{i}\to 0}},\qquad$ (26)
@@
Eq. (27) implies that $w\sim\sqrt{\lambda}=m^{-2}$, thus the scaling exponent of the field $\phi$, a.k.a. the roughness exponent $\zeta$, is
$\zeta=2,$ (28)
which also holds for the action (VII). By construction, the partition function of the latter over all bosonic and Grassmann fields is $1$,
@@
This function has a limit,
$\tilde{\Delta}_{\rm Susy}(w):=\lim_{m\to 0}\Delta_{\rm Susy}(wm^{-2},m^{-4}).$ (31)
It is non-trivial to show that the functions defined in Eqs. (31) and (24) agree (see the appendix and [23]),
@@
Thus what we obtained for the simple model (6) also applies to the disordered system defined by the action (VII).
@@ VIII Applications
Our results agree up to 1-loop order with that for disordered elastic manifolds in equilibrium and at depinning [17, 19, 18, 20, 21, 22]. Beyond that, amplitudes are different in the $\epsilon$-expansion, and there are additional anomalous terms which are hard to recuperate [23]. While our model can formally be derived from a field theory in equilibrium, we do not believe $\tilde{\Delta}_{\rm FRG}(w)$ to be relevant for a specific physical situation, even though the predicted roughness exponent is equal to that of depinning, and the shape of $\tilde{\Delta}_{\rm FRG}(w)$ on Fig. 3 is between a driven particle and Sinai’s model, both relevant in $d=0$.
Given these caveats, we turn to the strengths of our approach. Our model contains all ingredients of functional renormalization: it shows that the perturbative series is Borel summable, how the limit of strong coupling is reached, that it cannot be inferred from the large-order behavior, and how universality emerges. By connecting to a formulation via superfields, we establish the connection to field theory. The zero-dimensional limit is relevant for DNA unzipping [30] and RNA/DNA peeling [28], providing a concrete physical application. Since this limit retains all relevant physics, as e.g. avalanches, and can be assessed in an expansion in $\epsilon=4-d$, the model (7)-(8) is key in understanding functional RG and its $\epsilon$-expansion; as a solution of the model (1), even though somehow trivial, is key in understanding the $\epsilon$-expansion in $\phi^{4}$-theory.
Our work poses a solid framework for the strong-coupling behavior in functional renormalization, constraining the large-order and strong-coupling behavior in dimension $d>0$. We also saw that to define the path integral non-perturbatively, one needs to specify the integration contours, and constrain variables to part of their physically allowed domains. This restricts theories in dimension $d>0$.
@@ Appendix A Field Theory
Building on the Susy formulation of [38], Ref. [31] introduces two physical copies located at $u_{1}$ and $u_{2}$, which are subject to confining potentials displaced by $w$, s.t. their difference $\phi:=u_{1}-u_{2}$ has expectation $\left<\phi\right>=w$; its center of mass is $u:=(u_{1}+u_{2})/2$. The field theory, given in Eq. (36) of [31] reads
@@
After Eq. (20) we stated that the latter can be derived from Eq. (6) if $\tilde{\phi}\in i\mathbb{R}$ and $\phi>0$. We use the same prescription to pass from Eq. (34) to Eq. (35), checking that ${\cal Z}=1$ in Eq. (35). We then evaluated Eqs. (30)-(31) perturbatively and numerically, proving Eq. (32) (for details see [23]).
@@ Acknowledgements.
We are grateful to Andrei Fedorenko for stimulating discussions and many deep questions. We profited from exchanges with Costas Bachas and Edouard Brézin, and feedback from the referees.